I need help with this equation please--Supposed to solve the equation and check for extraneous solutions.
√(a-1)-5=1

Answer by profemmanuel2q@yahoo.com(15)

√(a-1)-5=1
√(a-1)= 1 +5
√(a-1)= 6 square both sides to get ride of the sqaure root
a - 1 = 36
a = 36 + 1
a = 37